OpenStudy (anonymous):

The legs of a right triangle are 4 units and 7 units long. To the nearest whole number, what is the length of the hypotenuse? 8 units 9 units ten units eleven units

geerky42 (geerky42):

Use pythagorean theorem? \[\Large a^2+b^2=c^2\]

geerky42 (geerky42):



OpenStudy (anonymous):

4^2 + 7^2 = c^2 16 + 49 = 65^2 sqrt of 65 = 8.06 units So c = 8 to the nearest whole number Therefore answer is 8 units